# Rapid Brightening of 3I/ATLAS Ahead of Perihelion

### v1

@@ Abstract
Draft version October 28, 2025
@@ 1 Introduction
Interstellar comet 3I/ATLAS was discovered on 2025 July 1 at a heliocentric distance au while inbound toward its 2025 October 29 perihelion at au (L. Denneau et al., 2025). At the time, it was comparable in brightness at similar to the previous interstellar comet, 2I/Borisov (D. Z. Seligman et al., 2025). However, that similarity turned out to be coincidence of timing, as 2I brightened slowly with decreasing heliocentric distance on its approach, with (observer distance -corrected) brightness scaling as with only —close to that of an inert, reflecting object—over –2 au. Meanwhile, 3I varied in its rate of brightening, but exhibited a much steeper overall over –2 au (Q. Ye et al., 2025; J. Tonry et al., 2025; D. Jewitt & J. Luu, 2025), perhaps reflecting a difference in nucleus properties or the faster speed at which it approached the Sun.
@@
Optical observations, however, rapidly became difficult from the ground at au as the comet approached its 2025 October 21 superior conjunction at only solar elongation from Earth, leaving a gap in this dataset just as the comet approaches its maximum heating at perihelion. However, the comet remained observable during this period to a number of spacecraft around the solar system that are not designed for comet observations, but can nonetheless help to fill in this gap (T. M. Eubanks et al., 2025). Among these spacecraft are several solar observatories with cameras that continually monitor the corona and inner heliosphere, but which also routinely record the passage of comets that happen to cross their fields of view (K. Battams & M. M. Knight, 2017). Here, we present observations from solar coronagraphic and heliospheric imagers onboard the STEREO-A, SOHO, and GOES-19 spacecraft, and as well as a cursory analysis of their implications for the 3I’s activity on its final approach toward perihelion.
@@ 2.1 Instruments
STEREO (“Solar TErrestrial RElations Observatory”; M. L. Kaiser et al., 2008) is comprised of two effectively identical spacecraft launched in 2006: STEREO-A, which orbits slightly interior to/more quickly than Earth, and STEREO-B, which orbits slightly exterior to/more slowly than Earth. STEREO-B was malfunctioned in and has been defunct since 2014 (D. A. Ossing et al., 2018), so only STEREO-A observed 3I. More specifically, the spacecraft carries a SECCHI (“Sun Earth Connection Coronal and Heliospheric Investigation”; R. A. Howard et al., 2008) instrument suite, of which we report observations from two cameras:
@@ 2.2 Data
We processed images from all instruments similarly following a procedure similar to that used by Q. Zhang et al. (2023), and started from the level-2 HI1, level-0 COR2, level-0.5 C3 (plus level-1 vignetting/bias correction), and level-1b CCOR-1 data. We first derived astrometric solutions for all coronagraph (COR2, C3, and CCOR-1) frames using Gaia DR3 (Gaia Collaboration et al., 2023), but used the existing HI1 solutions. We subtracted a stray light/corona model from all coronagraph frames, derived from an average of all frames in each dataset; level-2 HI1 data already includes this correction. We also subtracted a stellar background from each HI1 frame, derived from frames several hours away to avoid self-subtraction of the moving comet.
@@
Figure 1 shows the stack on the comet through each of the four instruments. The CCOR-1 stack resolves the comet as extended with a diameter coma, compared to a similarly stacked star nearby that serves as an approximate point spread function (PSF). No tail is clearly visible, although the duration of the stack combined with the low () phase angle means that an ion tail—which could vary in direction with the solar wind from the radial direction (Y. R. Fernández et al., 1997)—could theoretically point in any projected direction on the individual frames, so may simply be smeared out, perhaps contributing to the apparent coma. Any physically antisunward dust tail would also be highly foreshortened by this geometry. Note also that the sunward direction rotated by over the time range of the stack, so any angularly sunward-aligned structures would likewise be smeared out; however, the sunward-aligned stack does not appear qualitatively different from the presented north-aligned stack, so is not shown.
@@
For time series photometry, we stacked the cutouts over over 1 day windows, except for COR2 where we split the data into a pair of 2.5 day/1.7 day stacks with similar S/N, due to the low total S/N of that dataset being insufficient for useful 1 day stacks. We measured the flux within radii of for HI1, for COR2, and for both C3 and CCOR-1—which we estimate capture 90% of the comet’s flux visible above the noise level in the full image stack for each dataset—and converted them to the conventional solar magnitude system (i.e., Sun as at au, its Johnson magnitude; C. N. Willmer 2018), which is convenient for distinguishing differences from solar color. We used the existing photometric calibrations of S. Tappin et al. (2022) for HI1 and Q. Zhang et al. (2023) for C3. For COR2 and CCOR-1, we followed the same procedure Q. Zhang et al. (2023) used for LASCO, and derived zero-point magnitudes of 12.5 and 14.7 from 2025 April and May observations of the Sun-like star 39 Tau, respectively.
@@
Figure 2 presents all of the resulting photometry that detected the comet at , corrected to a common au. No phase corrections were applied, as the data cover a narrow phase angle range –, and the necessary dust-to-gas brightness ratio such a correction would depend on is unknown. Note, however, that any direct photometric fit across all points would implicitly assume the comet is solar-colored between the relevant instrument bandpasses, which is not necessarily true, as demonstrated by the offset between the contemporaneous CCOR-1 and LASCO C3 Clear photometry. While all four bandpasses have similar effective wavelengths of 600–700 nm, the variations in their spectral profiles still produce differences in sensitivity to common optical gas species, particularly CN, C2, and NH2 (Q.-Z. Ye et al., 2014; G. H. Jones et al., 2018).
@@
LASCO actually has the capability to directly measure color: while it takes the vast majority of its data through its Clear filter (including the data used for the time series photometry in Figure 2), it also presently takes one full resolution frame per day through each of its Blue, Orange, DeepRd (“deep red”), and IR (infrared) color filters, whose properties are detailed in G. Brueckner et al. (1995) and A. Thernisien et al. (2006). Stacking all the frames through each color filter yields detections of 3I in only the Blue and Orange stacks, in which the comet was respectively () mag and () mag brighter than in the equivalent Clear-filtered stack—distinctly bluer than solar color, in contrast to the comet’s red dust (C. Opitom et al., 2025; T. Kareta et al., 2025). The Blue bandpass efficiently transmits the C2 Swan bands, while the Orange bandpass is less sensitive to C2 but efficiently transmits several NH2 bands (H. Kawakita & J.-i. Watanabe, 2002; G. H. Jones et al., 2018), suggesting that the presence of these species could be at least partly responsible for the observed deviations from solar colors. The LASCO C3 photometry can also be compared with the contemporaneous CCOR-1 photometry, which are measured with the same photometric aperture size. CCOR-1’s bandpass is similar to, but 40 nm bluer than, the C3 Clear bandpass, increasing CCOR-1’s relative sensitivity to the C2 Swan bands. Indeed, the comet’s CCOR-1 brightness falls between the C3 Blue and Clear values, with a CCOR-1–Clear color of () mag. Table 2 provides the observed (i.e., not -corrected) magnitudes through each filter.
@@
Unlike C3 Clear and CCOR-1, neither HI1 nor COR2 are sensitive to the C2 Swan bands. However, both instruments have much narrower primary bandpasses covering multiple NH2 bands, offsetting the lack of C2 sensitivity with increased NH2 sensitivity. While these NH2 bands are normally much fainter than the C2 Swan bands (A. L. Cochran et al., 2015), 3I has been found to be C2-depleted (D. Hutsemékers et al., 2025), which may improve the trade balance if C2 is still depleted. HI1’s red and blue leaks also provide sensitivity to CN, while it remains unknown if COR2 has similar leaks. A rigorous quantitative analysis requires an accurate inventory of the comet’s optical emission and dust brightness, neither of which is available.
For our cursory analysis, we consider that the comet’s brightness in HI1 and COR2 would likely more similar to that in C3 Clear than in CCOR-1, and “correct” the CCOR-1 photometry to match the equivalent C3 Clear brightness (i.e., by applying a +0.4 mag offset) for the global fit. We then also allow for the possibility that HI1/COR2 brightness is offset from C3 Clear by a similar 0.4 mag. With these color corrections, we obtain (with a corresponding normalized magnitude at au)—much steeper than the previously reported for the comet’s earlier brightening trend at au (D. Jewitt & J. Luu, 2025).
@@ 3 Discussion
Condition 1 is plausible given the comet’s lack of a prominent dust tail and blue color, but significant dust contribution to the brightness cannot be ruled out with just the presented data—particularly for HI1 and COR2, which are insensitive to C2. NH2 is the major gas species that HI1 and C3 Clear (and CCOR-1, corrected to C3 Clear)—which anchor the two ends of the fit—are both sensitive to. Condition 2 is satisfied for NH2, as scale length for destruction (U. Fink et al., 1991) is contained within all the aperture radii we used, apart from the low S/N COR2 photometry that negligibly contributes to the fit. The Swings effect is usually neglected for NH2 when deriving their production rates (A. L. Cochran et al., 2015), satisfying condition 3. The status of condition 4 is not directly known, although the closeness of the in the comparison above and the possible satisfaction of the other three conditions hints at its possible validity.
@@ 4 Conclusions
The comet appears distinctly bluer than the Sun in LASCO/CCOR-1 color photometry—in contrast to earlier observations showing the comet’s dust to be red—suggesting that gas emission, likely from C2 and possibly NH2, contributes a sizable fraction of the overall visible brightness.

### v2

@@ Abstract
Publications of the Astronomical Society of the Pacific, in press
@@ 1 Introduction
Interstellar comet 3I/ATLAS was discovered on 2025 July 1 at a heliocentric distance au while inbound toward its 2025 October 29 perihelion at au (L. Denneau et al., 2025). At the time, it was comparable in brightness at similar to the previous interstellar comet, 2I/Borisov (D. Z. Seligman et al., 2025). However, that similarity turned out to be coincidence of timing, as 2I brightened slowly with decreasing heliocentric distance on its approach, with (observer distance -corrected) brightness scaling as with only —close to that of an inert, reflecting object—from the earliest pre-discovery observations at au to its perihelion at au. Meanwhile, 3I varied in its rate of brightening, but exhibited a much steeper overall from its earliest pre-discovery observations at through the pre-perihelion observations at au (Q. Ye et al., 2025; J. Tonry et al., 2025; D. Jewitt & J. Luu, 2025), perhaps reflecting a difference in nucleus properties or the faster speed at which it approached the Sun.
@@
Optical observations, however, rapidly became difficult from the ground at au as the comet approached its 2025 October 21 superior conjunction at only solar elongation from Earth, leaving a gap in this dataset just as the comet approaches its maximum heating at perihelion. However, the comet remained observable during this period to a number of spacecraft around the solar system that are not designed for comet observations, but can nonetheless help to fill in this gap (T. M. Eubanks et al., 2025). Among these spacecraft are several solar observatories with cameras that continually monitor the corona and inner heliosphere, but which also routinely record the passage of comets that happen to cross their fields of view (K. Battams & M. M. Knight, 2017). Here, we present observations from solar coronagraphic and heliospheric imagers onboard the STEREO-A, SOHO, and GOES-19 spacecraft, as well as a cursory analysis of their implications for the 3I’s activity on its final approach toward perihelion.
@@ 2.1 Instruments
STEREO (“Solar TErrestrial RElations Observatory”; M. L. Kaiser et al., 2008) is comprised of two effectively identical spacecraft launched in 2006: STEREO-A, which orbits slightly interior to/more quickly than Earth, and STEREO-B, which orbits slightly exterior to/more slowly than Earth. STEREO-B malfunctioned in, and has been defunct since, 2014 (D. A. Ossing et al., 2018), so only STEREO-A observed 3I. The spacecraft carries a SECCHI (“Sun Earth Connection Coronal and Heliospheric Investigation”; R. A. Howard et al., 2008) instrument suite, of which we report observations from two cameras:
@@ 2.2 Data
We processed images from all instruments similarly following a procedure similar to that used by Q. Zhang et al. (2023), and started from the level-2 HI1333Available online at https://stereo-ssc.nascom.nasa.gov/data/ins_data/secchi/secchi_hi/L2/a/img/hi_1/20250926/ for 2025 September 26 data; replace the 20250926 in the URL with the date of interest in YYYYMMDD format., level-0 COR2444Available online at https://stereo-ssc.nascom.nasa.gov/data/ins_data/secchi/L0/a/img/cor2/20251001/ for 2025 October 1 data; likewise, replace the 20251001 with the date of interest, level-0.5 C3555Available online at https://lasco-www.nrl.navy.mil/lz/level_05/251020/c3/ for 2025 October 20 data; replace the 251020 with the date of interest in YYMMDD format. (plus level-1 vignetting/bias correction), and level-1a CCOR-1 data666Available online at https://noaa-nesdis-swfo-ccor-1-pds.s3.amazonaws.com/index.html#SWFO/GOES-19/CCOR-1/ccor1-l1a/.. We first derived astrometric solutions for all coronagraph (COR2, C3, and CCOR-1) frames using Gaia DR3 (Gaia Collaboration et al., 2023), but used the existing HI1 solutions. We subtracted a stray light/corona model from all coronagraph frames, derived from an average of all frames in each dataset; level-2 HI1 data already includes this correction. We also subtracted a stellar background from each HI1 frame, derived from frames several hours away to avoid self-subtraction of the moving comet.
@@
Figure 1 shows the stack on the comet through each of the four instruments. The CCOR-1 stack resolves the comet as extended with a diameter coma, compared to a similarly stacked star nearby that serves as an approximate point spread function (PSF). No tail is clearly visible. Note, however, that while an ion tail nominally points in antisunward direction with a aberration lag at the comet’s 70 km s-1 near perihelion through a typical 400 km s-1 solar wind, it can vary from this direction by another with directional deviations in the solar wind (Y. R. Fernández et al., 1997). The long duration of the stack combined with the low () phase angle means that such an ion tail could theoretically point in any projected direction on the individual frames, and thus, may simply be smeared out in the stack, perhaps contributing to the apparent coma. In addition, any physically antisunward dust tail would also be highly foreshortened by this geometry. Note also that the sunward direction rotated by over the time range of the stack, so any angularly sunward-aligned structures would likewise be smeared out; however, the sunward-aligned stack does not appear qualitatively different from the presented north-aligned stack, so is not shown.
For time series photometry, we stacked the cutouts over over 1 day windows, except for COR2 where we split the data into a pair of 2.5 day/1.7 day stacks with similar S/N, due to the low total S/N of that dataset being insufficient for useful 1 day stacks. We measured the flux within radii of for HI1, for COR2, and for both C3 and CCOR-1. These apertures far exceed the dust coma/tail reported by D. Jewitt & J. Luu (2025) in ground-based observations contemporaneous with the HI1, and are also comparable to or larger than the Haser scale lengths for most optical gases each bandpass is sensitive to (A. L. Cochran et al., 1992), thus capturing nearly all of the comet’s expected dust and gas flux. We verified this conclusion by performing photometry on the full image stack for each dataset (which has higher S/N than individual time series stacks, permitting photometry with larger apertures) with both the adopted apertures and ones 2 the radii, and found the former captured 90% of the flux within the latter.
@@
We then converted the instrumental fluxes to the conventional solar magnitude system (i.e., Sun as at au, its Johnson magnitude; C. N. Willmer 2018), which is convenient for distinguishing differences from solar color. We used existing time-dependent photometric calibrations of S. Tappin et al. (2022) for HI1 and Q. Zhang et al. (2023) for C3, extrapolated to the appropriate epochs in 2025. For COR2 and CCOR-1, we followed the same procedure Q. Zhang et al. (2023) used for LASCO, and derived zero-point magnitudes of and from 2025 April and May observations of the solar analog star 39 Tau ( magnitude 5.9, differing from solar color by 0.1 mag over the relevant wavelengths; T. L. Farnham et al., 2000), respectively.
Figure 2 presents all of the resulting photometry that detected the comet at , corrected to a common au. No phase correction has been applied, as the data cover a narrow phase angle range – corresponding to only 0.3 mag variation in dust brightness under the standard Schleicher–Marcus phase function (D. G. Schleicher & A. N. Bair, 2011; J. N. Marcus, 2007), which is further diluted by an unknown dust-to-gas brightness ratio. Moreover, the phase function of even the dust alone may not necessarily match that of solar system comets given the unusual polarization of sunlight scattered by 3I’s dust that is indicative of unusual grain properties and scattering behavior (Z. Gray et al., 2025).
@@
Note, however, that any direct photometric fit across all points would implicitly assume the comet is solar-colored between the relevant instrument bandpasses, which is not necessarily true, as demonstrated by the offset between the contemporaneous CCOR-1 and LASCO C3 Clear photometry. While all four bandpasses have similar effective wavelengths of 600–700 nm, the variations in their spectral profiles still produce differences in sensitivity to common optical gas species, particularly C2, NH2, and CN (Q.-Z. Ye et al., 2014; G. H. Jones et al., 2018).
@@
LASCO actually has the capability to directly measure color: while it takes the vast majority of its data through its Clear filter (including the data used for the time series photometry in Figure 2), it also presently takes one full resolution frame per day through each of its Blue, Orange, DeepRd (“deep red”), and IR (“infrared”) color filters, whose properties are detailed in G. Brueckner et al. (1995) and A. Thernisien et al. (2006). Stacking all the frames through each color filter yields detections of 3I in only the Blue and Orange stacks. The corresponding photometry yields a Blue–Clear color of () mag and an Orange–Clear color of () mag—distinctly bluer than solar color, in contrast to the comet’s red dust (C. Opitom et al., 2025; T. Kareta et al., 2025). The Blue bandpass efficiently transmits the C2 Swan bands, while the Orange bandpass is less sensitive to C2 but efficiently transmits several NH2 bands (H. Kawakita & J.-i. Watanabe, 2002; G. H. Jones et al., 2018), suggesting that the presence of these species could be at least partly responsible for the observed deviations from solar colors. The LASCO C3 photometry can also be compared with the contemporaneous CCOR-1 photometry, which are measured with the same photometric aperture size from nearly the same location (GOES-19, in geosynchronous orbit, was only 0.4% farther from the comet than SOHO, in its L1 halo orbit). CCOR-1’s bandpass is similar to, but 40 nm bluer than, the C3 Clear bandpass. The difference puts the strongest, band of C2—which falls just blueward of C3 Clear’s FWHM interval—cleanly within CCOR-1’s bandpass, substantially elevating its sensitivity to C2 (K. Sivaraman et al., 1988). Indeed, the comet’s CCOR-1 brightness falls between the C3 Blue and Clear values, with a CCOR-1–Clear color of () mag. Table 2 provides the observed (i.e., not -corrected) magnitudes through each filter from which these colors were derived.
@@
Unlike C3 Clear and CCOR-1, neither HI1 nor COR2 are sensitive to the C2 Swan bands. However, both instruments have much narrower primary bandpasses covering multiple NH2 bands, offsetting the lack of C2 sensitivity with increased NH2 sensitivity. While these NH2 bands are normally much fainter than the C2 Swan bands (A. L. Cochran et al., 1992), 3I has been found to be C2-depleted (D. Hutsemékers et al., 2025), which may improve the trade balance if C2 is still depleted. HI1’s blue leak also provides slight sensitivity to CN and C3, while it remains unknown if COR2 has similar leaks. Figure 3 provides a visualization of these bandpasses relative to these and optical gas emission features of a typical solar system comet. A rigorous quantitative analysis requires an accurate inventory of the comet’s optical emission and dust brightness, neither of which is available.
For our cursory analysis, we consider that the comet’s brightness in HI1 and COR2 would likely be more similar to that in C3 Clear than CCOR-1 given C3 Clear’s much lower relative C2 sensitivity, and therefore convert the CCOR-1 photometry to provide the equivalent C3 Clear brightness (i.e., by applying a +0.4 mag offset) for the global fit. Given that the C2 Swan band largely captured by CCOR-1, but excluded by C3 Clear, comprises most of the brightness of C2 (K. Sivaraman et al., 1988), that C2 sensitivity is the primary difference between C3 Clear and HI1/COR2, and that any gas emission in HI1/COR2 is either also within the C3 Clear bandpass or only weakly transmitted in a red or blue leak, we expect the offset between HI1/COR2 and C3 Clear will likely be no more than the CCOR-1–Clear difference. Therefore, we incorporate into our global fit a potential offset of HI1/COR2 brightness from C3 Clear by 0.4 mag using Monte Carlo sampling to fit the light curve with a corresponding ensemble of HI1/COR2 offsets. With these color allowances, we obtain (with a corresponding normalized C3 Clear magnitude at au)—much steeper than the previously reported for the comet’s earlier brightening trend at au (D. Jewitt & J. Luu, 2025).
@@ 3 Discussion
Condition 1 is plausible given the comet’s lack of a prominent dust tail and blue color, but a dominant dust contribution to the brightness remains compatible with the presented data—particularly for HI1 and COR2, which are insensitive to C2. NH2 is the major gas species that HI1 and C3 Clear (and CCOR-1, corrected to C3 Clear)—which anchor the two ends of the fit—are both sensitive to. Condition 2 is satisfied for NH2, as its Haser scale lengths (U. Fink et al., 1991) are contained within all the aperture radii we used, apart from the low S/N COR2 photometry that negligibly contributes to the fit. The Swings effect is usually neglected for NH2 when deriving their production rates (A. L. Cochran et al., 1992), satisfying condition 3. The status of condition 4 is not directly known, although the closeness of the in the comparison above and the possible satisfaction of the other three conditions hints at its possible validity.
@@ 4 Conclusions
The comet appears distinctly bluer than the Sun in LASCO/CCOR-1 color photometry—in contrast to earlier observations showing the comet’s dust to be red—suggesting that gas emission, likely from C2 (and possibly NH2), contributes a sizable fraction of the overall visible brightness.